# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

### v3

@@ 3.2 Synthetic Training Data Generation
3.2.1 Data Generation with Linear Factor Models
We generate data using a simple linear factor model (LFMs) (linearfactor). LFMs are commonly used in matrix completion literature to prove error bounds for matrix completion algorithms (netflixprize; rechtexact). This family of models generates a data matrix $Y\in\mathbb{R}^{m\times n}$ by assuming the data lies on or near a low-dimensional subspace. The simplest case generates the data matrix $Y$ as the inner product of two lower-rank latent factor matrices, $U\in\mathbb{R}^{m\times k}$ and $V\in\mathbb{R}^{n\times k}$, where $k\ll n,m$ is the rank:
$Y=UV^{T}.$ (4)
To generate diverse datasets, the latent vectors (rows of $U$ and $V$) are sampled from a variety of distributions, including Gaussian, Laplace, Student’s t, spike-and-slab (a mixture of a Dirac delta at zero and a Gaussian), and Dirichlet.
When training, we experimented with several classes of data-generating processes (DGP), including matrices from nonlinear factor models and structural causal models (SCM) similar to the ones used in TabPFN. We found that training on SCM’s proved too computationally expensive, but were able to train a model on nonlinear factor models. We found that the model trained on linear factor models performed the best, as shown in Tab. 9 in Sec. A.3. We leave it as future work to explore other DGPs which might further enhance the underlying model.
3.2.2 Missingness Patterns for Training and Evaluation
After generating a complete data matrix $X^{*}\in\mathbb{R}^{m\times n}$, we introduce missingness by applying a masking matrix $M\in\{0,1\}^{m\times n}$, where the entry value $M_{ij}=1$ if and only if $(i,j)\in\Omega_{\mathrm{obs}}$. To ensure TabImpute is robust and generalizable to the variety of ways data can be missing in real-world scenarios, we pre-train on a comprehensive stock of synthetic datasets with several missingness patterns. For convenience, we define $p_{ij}=\mathbb{P}(M_{ij}=1)$, the propensity of each entry in $X^{*}$.
We include $13$ different missingness patterns: $1$ MCAR, $1$ MAR pattern, and $11$ MNAR patterns. For examples of these, see Fig. 2. MNAR patterns often stump standard imputation methods and yet are extremely common in the real world.
@@ 3.2.2 Missingness Patterns for Training and Evaluation
For MAR missingness, the probability of an entry being missing depends only on the observed values $X$. In other words, the randomness in MAR can be explained by conditioning on observed factors. Additionally, every entry has a positive probability of being observed (i.e., $p_{ij}>0$). We simulate MAR through column-wise MAR, denoted Col-MAR: we choose several columns as predictor columns and use those values to mask entries in other columns. This is similar to the MAR approach taken in jarrett2022hyperimpute.
MNAR Patterns:
For the most complex missingness class, MNAR, the probability of an entry being missing can depend on unobserved factors. Note that MNAR patterns are significantly more difficult to handle systematically, often requiring specialized methods for a specific kind of MNAR pattern (van2012flexible). Due to the flexibility of our entry-wise featurization, TabImpute can produce imputations for these highly complex scenarios, even when columns are completely missing, such as in panel-data missingness patterns. HyperImpute was tested on two MNAR patterns in the Appendix of jarrett2022hyperimpute, one where values are further masked after an MAR pattern and another where values outside a certain range are masked. We build on this work by testing on $11$ MNAR patterns (see Sec. A.2 and Tab. 10 for details). We implement a range of MNAR patterns to simulate plausible real-world scenarios. For example, we utilize the expressiveness of neural networks to create random propensity functions MNAR patterns, use bandit algorithms to induce column-adaptive missing patterns, simulate panel data missingness where some features are removed later, censoring where sensor readings fall outside a detectable range, and survey data artifacts like respondent polarization and skip-logic.
@@ 3.3 Training TabImpute
We train our model to predict unobserved values under several missingness patterns simultaneously. For the final TabImpute model, we trained only on MCAR missingness because we found our model generalized to the other patterns well without including them explicitly in training. In fact, we found that including other missingness patterns degraded overall performance. We discuss these results further below. We use the prior-data fitted negative log likelihood (NLL) loss proposed in muller2022transformers. Like other PFNs with continuous numerical output, we use the Riemann distribution output also proposed in muller2022transformers. Since we can generate an unlimited amount of synthetic data, we do not reuse any synthetic data and only do one gradient pass per batch of datasets. This allows our model to learn the underlying data-generating process and missingness mechanisms without risk of memorization. We use a learning rate of $0.0001$, a batch size of $16$, and train on around $1.9$ million synthetic datasets.
Remark: We had initially attempted to train the model sequentially on one missingness pattern at a time, but found that the network always experienced catastrophic forgetting (mccloskey1989catastrophic) irrespective of learning rate (i.e., it forgot how to handle the previous missingness patterns). After this, we attempted to mix several missingness types together: MCAR, MAR, and Self-Masking-MNAR. However, we found that the model’s performance degraded overall because it focused only Self-Masking-MNAR. We attempted several mitigation techniques including pattern weight schedules and GradNorm chen2018gradnorm, a gradient normalization technique for multi-task learning, but found them unsuccessful in improving performance.
@@ 4 Results on OpenML Datasets: MissBench
To evaluate TabImpute against other methods, we introduce MissBench: a missing-data imputation benchmark using $42$ OpenML (OpenML2013) tabular datasets with $13$ synthetic missingness patterns. For every dataset and missingness pattern, we test each method’s ability to impute masked values. The $42$ OpenML datasets span domains such as medicine, engineering, and education. The missingness patterns include $1$ MCAR pattern, $1$ MAR pattern, and $11$ MNAR patterns. We provide details for the MNAR patterns in Sec. A.2.
We test a suite of imputation methods on MissBench: column-mean imputation (hawthorne2005imputing), SoftImpute (hastie2015matrix), MissForest (missforest2011), iterative chained estimators (ICE/MICE) (ice2011; royston2011multiple), GAIN (gain2018), MIWAE (mattei2019miwae), an optimal transport-based method (muzellec2020missing), and TabPFN’s imputation method (see Sec. 1). We also test imputing categorical variables via one-hot encoding in Sec. A.1 and report $R^{2}$ values in Tab. 8.
@@
To ensure a fair comparison across datasets with different scales and inherent difficulties, we report a normalized accuracy score. In particular, for each dataset and missingness pattern, we first calculate the standard Root Mean Squared Error (RMSE) for every imputation method as $\big(\frac{1}{|\Omega|}\sum_{(i,j)\in\Omega}\big(X^{\text{true}}_{ij}-X^{\text{imputed}}_{ij}\big)^{2}\big)^{1/2}$}, where $\Omega$ denotes the set of missing entries. We then perform a min-max normalization on these RMSE scores across all methods for that specific task:
$\text{Normalized RMSE}=\frac{\text{RMSE}_{\text{method}}-\min(\text{RMSE}_{\text{all methods}})}{\text{max}(\text{RMSE}_{\text{all methods}})-\min(\text{RMSE}_{\text{all methods}})}$ (5)
This normalization maps the best-performing method to 0 and the worst to 1. Finally, we define Imputation Accuracy as 1Normalized RMSE1-$\text${Normalized RMSE}, where higher values indicate better performance.
The final scores we report represent imputation accuracy averaged across all $42$ datasets and $13$ missingness patterns. Dataset sizes range from $50\times 5$ to $170\times 55$. We evaluate only on datasets with numerical values without pre-existing missingness before applying synthetic patterns. Specific datasets are listed in Tab. 11 (Sec. A.3).
@@
Tab. 1 presents results for each missingness pattern as well as overall performance. TabImpute achieves the best overall performance and for nearly all individual patterns. For completeness, we list the performance of methods not shown in the table in Tab. 6 and non-normalized RMSE values in Tab. 12. TabImpute performs best under high missingness conditions (Fig. 4, Sec. A.1), which is expected since it leverages generative pre-training rather than relying solely on available dataset information like discriminative methods. We also include column-wise 1-Wasserstein distance comparisons in Tab. 7 in Sec. A.3. In terms of the 1-Wasserstein distance, TabImpute and EWF-TabPFN do not perform as well as prior methods; however, we leave this for future work to investigate further.
We train on the pattern above the dashed line. The number of samples used to calculate standard error are the number of datasets: $42$.
Pattern TabImpute EWF-TabPFN HyperImpute MissForest
MCAR 0.882 ± 0.021 0.817 ± 0.024 0.804 ± 0.031 0.867 ± 0.022
NN-MNAR 0.900 ± 0.019 0.876 ± 0.020 0.760 ± 0.036 0.818 ± 0.029
Self-Masking-MNAR 0.703 ± 0.038 0.682 ± 0.042 0.689 ± 0.039 0.638 ± 0.040
Col-MAR 0.877 ± 0.030 0.855 ± 0.027 0.814 ± 0.040 0.773 ± 0.039
Block-MNAR 0.890 ± 0.026 0.905 ± 0.026 0.873 ± 0.027 0.860 ± 0.023
Seq-MNAR 0.920 ± 0.014 0.905 ± 0.015 0.864 ± 0.030 0.829 ± 0.031
Panel-MNAR 0.915 ± 0.025 0.711 ± 0.048 0.865 ± 0.033 0.912 ± 0.020
Polarization-MNAR 0.804 ± 0.024 0.879 ± 0.021 0.622 ± 0.039 0.567 ± 0.033
Soft-Polarization-MNAR 0.759 ± 0.045 0.864 ± 0.024 0.711 ± 0.034 0.667 ± 0.041
Latent-Factor-MNAR 0.891 ± 0.021 0.883 ± 0.018 0.775 ± 0.036 0.834 ± 0.023
Cluster-MNAR 0.903 ± 0.014 0.871 ± 0.021 0.839 ± 0.028 0.830 ± 0.021
Two-Phase-MNAR 0.886 ± 0.026 0.851 ± 0.029 0.865 ± 0.031 0.873 ± 0.020
Censoring-MNAR 0.710 ± 0.032 0.593 ± 0.045 0.786 ± 0.041 0.671 ± 0.038
Overall 0.849 ± 0.008 0.823 ± 0.009 0.790 ± 0.010 0.780 ± 0.009
@@
In multiple imputation, a dataset is imputed multiple times independently. These datasets can then be used to provide better confidence intervals of downstream quantities van2012flexible. Given two imputation methods that have the same imputation accuracy, one would prefer a method with higher uncertainty to not provide false precision on imputed values van2012flexible; li2015multiple; jolicoeur2024generating. We assess performance for multiple imputation by repeating imputation 5 times per dataset and calculating the median absolute deviation (MAD), minimimum RMSE, and average RMSE, shown in Tab. 2. While most imputation methods require another full run in order to create another imputation, TabImpute and other PFN-based methods output a distribution. This distribution is then sampled 5 times to create 5 different imputations, following the methodology of ForestDiffusion jolicoeur2024generating. Thus, multiple imputation for TabImpute uses the same runtime as a single imputation.
TabImpute provides better diversity than methods of comparable imputation accuracy (MissForest and HyperImpute). The poor precision of MissForest and HyperImpute is shown through their worse minimum RMSE values. Note that in Tab. 2, TabImpute has a higher avergage RMSE value then MissForest and HyperImpute. This is because it has a higher diversity in its estimate (the samples have a larger spread). When comparing just the median estimate, though, TabImpute has a better accuracy, as shown in the first row in Tab. 1. Thus, TabImpute provides practitioners with the best of both worlds: If they require multiple imputation for uncertainty estimates, TabImpute provides diverse estimates, and if they require a single good imputation, TabImpute provides the most accurate estimate.
Table 2: Multiple Imputation Metrics: Dataset-level MAD and RMSE Metrics ± Standard Error
per Imputed Entry Across 5 Samples (MCAR, p=0.4).
Method MAD ($\uparrow$) Min RMSE ($\downarrow$) Avg RMSE ($\downarrow$)
@@
After training TabImpute on the MCAR pattern, we attempted to improve it by mixing in Col-MAR and Self-Masking-MNAR in the pre-training process. However, we found that the model consistently collapsed to only learning Self-Masking-MNAR, degrading performance on other patterns. This new model, though, performed far better on Self-Masking-MNAR than any other method, demonstrating the power of this architecture and pre-training to learn any pattern. We show the relative performance of three models trained on MCAR, Col-MAR, and Self-Masking-MNAR, respectively, in Tab. 3.
The model trained on Self-Masking-MNAR performs much better than the other models on this pattern. However, it does not generalize well to other patterns because the masks generated by the Self-Masking-MNAR pattern do not cover the types of masks generated by other patterns. On the other hand, when generating MCAR patterns millions of times, a small portion of these masks will look like Self-Masking-MNAR masks. Interestingly, the MCAR-trained model performed slightly better on Col-MAR patterns than the model trained on Col-MAR exclusively. One explanation for this is that the masks generated under MAR pattern are covered well by MCAR masks and MCAR masks produce a better underlying model. Note that the imputation accuracy scores are different for Tab. 1 and Tab. 3 because the MAR and Self-Masking-MNAR models are not included in the normalization in Tab. 1. However, the order of the imputation accuracy remains consistent.
Remark: We implemented an ensembling approach which weighted specialized model outputs by how well they predicted the observed values within a matrix, but found this approach to always favor the MCAR-trained model. We leave it for future work to determine a better method for ensembling or routing the specialized model predictions to better take advantage of specialized models.
Table 3: Imputation Accuracy ± Standard Error by Missingness Pattern for Specialized Models

### v4

@@ 3.1 Synthetic Training Data Generation
3.1.1 Data Generation: Linear Factor Models
We generate synthetic data using a linear factor model (LFM) (linearfactor). LFMs are commonly used in matrix completion literature to prove error bounds for matrix completion algorithms (netflixprize; rechtexact). This family of models generates a data matrix $Y\in\mathbb{R}^{m\times n}$ by assuming the data lies on or near a low-dimensional subspace. The simplest case generates the data matrix $Y$ as the inner product of two lower-rank latent factor matrices, $U\in\mathbb{R}^{m\times k}$ and $V\in\mathbb{R}^{n\times k}$, where $k\ll n,m$ is the rank: $Y=UV^{T}$. The latent vectors (rows of $U$ and $V$) are sampled from a variety of distributions including Gaussian, Laplace, and Student’s-t.
During pre-training, we experimented with several classes of data-generating processes (DGPs), including matrices from nonlinear factor models and structural causal models (SCMs) similar to those in TabPFN. Training on SCM’s proved too computationally expensive, but we were able to train a model on nonlinear factor models. The model trained on linear factor models performed the best, as shown in Tab. 9 in Sec. A.4. We leave it as future work to explain why LFMs work so well in the small-table regime.
3.1.2 Missingness Patterns
After generating a complete data matrix $X^{*}\in\mathbb{R}^{m\times n}$, we introduce missingness by applying a mask $M\in\{0,1\}^{m\times n}$, where the entry value $M_{ij}=1$ if $(i,j)\in\Omega_{\mathrm{obs}}$. To ensure TabImpute is robust and generalizable to the variety of ways data can be missing in real-world scenarios, we pre-train on millions of synthetic datasets, each with a different mask. For convenience, we define $p_{ij}=\mathbb{P}(M_{ij}=1)$, the propensity of each entry. We evaluate TabImpute on $13$ different missingness patterns: MCAR, $1$ MAR pattern, and $11$ MNAR patterns. For examples of these, see Fig. 2.
@@ 3.1.2 Missingness Patterns
For MAR missingness, the probability of an entry being missing depends only on the observed values $X$. In other words, the randomness in MAR can be explained by conditioning on observed factors. Additionally, every entry has a positive probability of being observed (i.e., $p_{ij}>0$). We simulate MAR through column-wise MAR, denoted Col-MAR: we choose several columns as predictor columns and use those values to mask entries in other columns. This is similar to the MAR approach taken in (jarrett2022hyperimpute).
MNAR:
For the most complex missingness class, MNAR, the probability of an entry being missing can depend on unobserved factors. Note that MNAR patterns are significantly more difficult to handle systematically, often requiring specialized methods for a specific kind of MNAR pattern (van2012flexible). Due to the flexibility of our entry-wise featurization, TabImpute can produce imputations for these highly complex scenarios. HyperImpute was tested on two MNAR patterns in the Appendix of (jarrett2022hyperimpute), one where values are further masked after an MAR pattern and another where values outside a certain range are masked. We build on this work by testing on $11$ MNAR patterns (see Sec. A.2 and Tab. 10 for details). We implement a range of MNAR patterns to simulate plausible real-world scenarios. For example, we utilize the expressiveness of neural networks to create random propensity functions MNAR patterns, use bandit algorithms to induce column-adaptive missing patterns, simulate panel data missingness where some features are removed later, and censoring where sensor readings fall outside a detectable range.
@@ 3.2 Training TabImpute
We use the negative log likelihood (NLL) loss proposed in (muller2022transformers), and like other PFNs with continuous numerical output, we use the Riemann distribution output also proposed in (muller2022transformers). Since we can generate an unlimited amount of synthetic data, we do not reuse any synthetic data by using one gradient pass per batch of datasets. This allows our model to learn the underlying data-generating process and missingness mechanisms without risk of memorization. We use the AdamW optimizer (loshchilov2017fixing), a learning rate of 1e-4, a batch size of $16$, and around $1.9$ million synthetic datasets.
We trained TabImpute only on MCAR missingness, finding that this model generalized well to the other patterns without explicitly including them in training. In fact, we found that including other missingness patterns degraded overall performance. We had initially attempted to train the model sequentially on one missingness pattern at a time, but found that the network always experienced catastrophic forgetting (mccloskey1989catastrophic) irrespective of learning rate (i.e., it forgot how to handle the previous missingness patterns). After this, we attempted to mix several missingness types together: MCAR, MAR, and Self-MNAR. However, the model’s performance degraded overall because it focused only Self-MNAR. We attempted several techniques, including weighting schedules and GradNorm (chen2018gradnorm), a gradient normalization technique for multi-task learning, but found them unsuccessful in improving performance.
Why MCAR Pre-Training Generalizes Beyond MCAR.
By generating millions of synthetic datasets under MCAR, the model is exposed to a broad and diverse set of missingness masks. While MNAR mechanisms induce biased mask distributions that emphasize specific structural patterns, many such masks overlap with the support of masks encountered under MCAR generation, just with different frequencies. Importantly, TabImpute does not model the underlying missingness dynamics, but rather the conditional data distribution given the observed data, $P(X_{ij}|X_{\Omega_{\text{obs}}})$. This design is well suited for a general-purpose imputation method that maximizes predictive accuracy across many missingness patterns without pattern-specific tuning. If the missingness mechanism is known a priori, though, specialized models perform much better, as shown in Tab. 3. Notably, our pipeline enables practitioners to easily train specialized models tailored to their missingness mechanisms without modifying the imputation algorithm.
The number of samples used to calculate standard error are the number of datasets: $42$.
Pattern TabImpute OT HyperImpute MissForest K-Nearest Neighbors
MCAR 0.904 ± 0.018 0.866 ± 0.019 0.814 ± 0.027 0.854 ± 0.025 0.772 ± 0.022
NN-MNAR 0.943 ± 0.014 0.873 ± 0.014 0.785 ± 0.035 0.838 ± 0.026 0.800 ± 0.020
Self-MNAR 0.775 ± 0.030 0.693 ± 0.033 0.751 ± 0.034 0.700 ± 0.034 0.680 ± 0.037
Col-MAR 0.921 ± 0.018 0.788 ± 0.032 0.868 ± 0.029 0.830 ± 0.025 0.872 ± 0.017
Block-MNAR 0.937 ± 0.013 0.857 ± 0.027 0.854 ± 0.027 0.850 ± 0.024 0.893 ± 0.014
Seq-MNAR 0.930 ± 0.015 0.853 ± 0.029 0.837 ± 0.031 0.802 ± 0.033 0.851 ± 0.027
Panel-MNAR 0.955 ± 0.008 0.918 ± 0.011 0.883 ± 0.031 0.897 ± 0.019 0.940 ± 0.011
Polarization-MNAR 0.825 ± 0.025 0.787 ± 0.021 0.631 ± 0.039 0.580 ± 0.033 0.658 ± 0.028
Soft-Polarization-MNAR 0.805 ± 0.047 0.773 ± 0.029 0.669 ± 0.036 0.625 ± 0.039 0.635 ± 0.035
Latent-Factor-MNAR 0.923 ± 0.016 0.870 ± 0.019 0.783 ± 0.038 0.835 ± 0.026 0.810 ± 0.022
Cluster-MNAR 0.933 ± 0.011 0.868 ± 0.016 0.844 ± 0.029 0.848 ± 0.021 0.807 ± 0.018
Two-Phase-MNAR 0.914 ± 0.012 0.822 ± 0.025 0.880 ± 0.029 0.881 ± 0.016 0.905 ± 0.015
Censoring-MNAR 0.766 ± 0.030 0.739 ± 0.021 0.810 ± 0.028 0.749 ± 0.023 0.561 ± 0.045
Overall 0.887 ± 0.007 0.824 ± 0.007 0.801 ± 0.009 0.791 ± 0.008 0.783 ± 0.009
@@ 4 Results on OpenML Datasets: MissBench
To evaluate TabImpute against other methods, we introduce MissBench: a missing-data imputation benchmark using $42$ OpenML (OpenML2013) small tabular datasets with $13$ synthetic missingness patterns. For every dataset and missingness pattern, we test each method’s ability to impute masked values. The $42$ OpenML datasets span domains such as medicine, engineering, and education. The missingness patterns include MCAR, Col-MAR, and $11$ different MNAR patterns, with details provided in Sec. A.2. We test a suite of imputation methods on MissBench: SoftImpute (hastie2015matrix), MissForest (missforest2011), iterative chained estimators (ICE/MICE) (ice2011; royston2011multiple), GAIN (gain2018), MIWAE (mattei2019miwae), an optimal transport-based method (muzellec2020missing), ForestDiffusion (jolicoeur2024generating), ReMasker (du2023remasker), CACTI (gorla2025cacti), and Col-TabPFN (see Sec. 1).
We report results across all $42$ datasets and $13$ missingness patterns. Dataset sizes range from $50\times 5$ to $170\times 55$. We evaluate only on datasets with numerical values without pre-existing missingness before applying synthetic patterns. Specific datasets are listed in Tab. 11 in Sec. A.4.
@@
To ensure a fair comparison across datasets with different scales and inherent difficulties, we report an accuracy score normalized across datasets. For each dataset and missingness pattern, we first calculate the root mean squared error (RMSE) (standardized per column by mean and standard deviation) for every imputation method as $\big(\frac{1}{|\Omega|}\sum_{(i,j)\in\Omega}\big(X^{\text{true}}_{ij}-X^{\text{imputed}}_{ij}\big)^{2}\big)^{1/2}$}, where $\Omega$ denotes the set of missing entries. We then perform a min-max normalization on these RMSE scores across all methods within each dataset:
$\text{Normalized RMSE}=\frac{\text{RMSE}_{\text{method}}-\min(\text{RMSE}_{\text{all}})}{\text{max}(\text{RMSE}_{\text{all}})-\min(\text{RMSE}_{\text{all}})}.$ (2)
Finally, following the normalized negative RMSE metric used in (hollmann2025accurate), we define Imputation Accuracy as 1Normalized RMSE1-$\text${Normalized RMSE}, so that higher values indicate better performance, analogous to accuracy metrics where larger values correspond to improved predictions.
@@
Tab. 1 presents results for each missingness pattern as well as overall performance. TabImpute achieves the best overall performance for nearly all individual patterns. For completeness, we list the performance of methods not shown in the table in Tab. 6 and non-normalized RMSE values in Tab. 12. TabImpute performs best under high missingness conditions (Fig. 5, Sec. A.1), which is expected since it leverages generative pre-training rather than relying solely on available dataset information like discriminative methods. In the Appendix, we report 1-Wasserstein distances in Tab. 7, report column-wise $R^{2}$ in Tab. 8, and test imputing categorical variables via one-hot encoding in Tab. 4.
For the other transformer-based methods, ReMasker (du2023remasker) and CACTI (gorla2025cacti), we observe lower relative performance than is reported in their papers. This difference is likely attributable to a regime mismatch: both methods are designed to operate on much larger tables. For example, Fig. 9 in (du2023remasker) shows ReMasker’s performance drops as dataset size decreases, eventually being overtaken by methods such as HyperImpute. Additionally, ReMasker is only tested on datasets with at least 308 rows and most having over 1,000 rows (Tab. 6 in (du2023remasker), and CACTI is only tested on datasets with at least 3,539 rows and half having over 15,000 rows (Tab. A9 in (gorla2025cacti)). In MissBench, however, we focus on small tables with at most 170 rows, resulting in a low-information setting. TabImpute is pre-trained to perform well with less information by leveraging its strong prior. These results highlight the difficulty of the small-table imputation problem and suggest that different imputation methods may be complementary across data regimes.
Discussion on Imputation Accuracy Metric.
Choosing an imputation metric is challenging because: (i) columns can have very different scales, (ii) datasets can have very different scales, and (iii) some datasets are inherently more difficult to impute than others (see Fig. 2 in (jarrett2022hyperimpute)). As a result, there is no consensus in the literature on the preferred method for measuring imputation accuracy. For example, MissForest reports RMSE standardized by the mean and standard deviation of the entire table’s missing values (missforest_python). HyperImpute reports non-standardized RMSE and column-wise 1-Wasserstein distance (jarrett2022hyperimpute). CACTI reports RMSE, 1-Wasserstein distance, and column-wise $R^{2}$ (gorla2025cacti). In our evaluation, we standardize RMSE values at the column level, since columns can differ substantially in scale (e.g., income versus age).
Further, a subtle issue that can arise in imputation benchmarking is data leakage through normalization. In particular, the evaluation pipelines of ReMasker (du2023remasker) and HyperImpute (jarrett2022hyperimpute) apply min–max scaling to each column before inducing missingness. So, if an extreme value is masked, the normalized value itself implicitly reveals information about the missing entry. To avoid data leakage, we standardize each column after inducing missingness.
@@
In multiple imputation, a dataset is imputed multiple times independently to provide better confidence intervals of downstream quantities (van2012flexible). Given two imputation methods that have the same imputation accuracy, one would prefer a method with higher uncertainty to not provide false precision on imputed values (van2012flexible; li2015multiple; jolicoeur2024generating). We assess performance for multiple imputation by repeating imputation 5 times per entry and calculating the median absolute deviation (MAD), minimum RMSE, and average RMSE, shown in Tab. 2 across the 5 samples per entry, following the methodology of ForestDiffusion (jolicoeur2024generating). While most methods require a full run to produce a new imputation, TabImpute outputs a distribution that can then be sampled from directly. Thus, multiple imputation for TabImpute runs as fast as single imputation.
TabImpute provides better diversity than methods of comparable imputation accuracy (MissForest and HyperImpute). The false precision of MissForest and HyperImpute is evident in their lower minimum RMSE values. Note that in Tab. 2, TabImpute has a higher average RMSE value than MissForest and HyperImpute. This is because TabImpute has greater diversity in its estimates (the samples have a larger spread). When comparing only the median-sample estimate, though, TabImpute has better accuracy, as shown in the first row of Tab. 1. Thus, TabImpute provides practitioners with the best of both worlds: If they require multiple imputation for uncertainty estimates, TabImpute provides diverse estimates, and if they require a single good imputation, TabImpute provides the most accurate estimate.
Table 2: Multiple Imputation Metrics: Per-Entry Metrics With 5 Samples Per Entry (MCAR, Probability of Missing = 0.4).
Method MAD Min. RMSE Avg. RMSE
@@
Here, we show the relative performance of three models trained on MCAR, Col-MAR, and Self-MNAR, respectively (see Tab. 3). The model trained on Self-MNAR performs much better than the other models on Self-MNAR, which is to be expected. However, it does not generalize well to other patterns because the masks generated by the Self-MNAR pattern do not cover the types of masks generated by other patterns. Interestingly, the MCAR-trained model performed slightly better on Col-MAR patterns than the model trained on Col-MAR exclusively. One possible hypothesis is that the masks generated under MAR pattern are well covered by MCAR masks, and MCAR masks produce a better underlying model. Note that the imputation accuracy scores are different for Tab. 1 and Tab. 3 because the MAR and Self-MNAR models are not included in the normalization in Tab. 1.
Table 3: Imputation Accuracy ± Standard Error by Missingness Pattern for Specialized TabImpute Models